26. Using the binomial distribution the probability that a sample of n = 10 from a process with proportion defective p = 0.1 will gives exactly one defective item is:

a. 0.1 b. 0.3874 c. 0.7361 d. 0.3588
Answer #1

p = 0.1 => q = 1-p = 0.9; r= 9 and n= 10

P(Defective =1) = nCr*pn-r*qr = 10C9*(0.1)*(0.9)9= 0.3874